I am studing the substitution Method.. and for some reason i can not seem to figure out how to do it. My teacher has been trying to work with me, but yet when I look at the problem I go blank on what to do.
I have an equation y = 2x + 7
y = 10x - 1 I dont understand how to do this...

Answer by xcentaur(357)

y = 2x + 7 .... [1]
y = 10x - 1 .... [2]


y=2x+7
2x=y-7
x=(y-7)/2 ...... [3]


Substituting value of [3] in [2],
y=10x-1
y=(10(y-7)/2)-1
y=5(y-7)-1
y=5y-35-1
y-5y=-35-1
-4y=-36
y=-36/-4
y=9


Also,putting y=9 in equation [1] we get,
y=2x+7
9=2x+7
9-7=2x
2x=2
x=1


ANS:
x=1,y=9


Hope this helps,
best of luck.